# WavLLM: Towards Robust and Adaptive Speech Large Language Model

### v1

@@ 1 Introduction
Amidst the broadening scope of abilities, speech stands out as a crucial form of human communication, prompting extensive research to equip large language models (LLMs) with speech perception capabilities (Shu et al., 2023; Wu et al., 2023; Tang et al., 2023; Chu et al., 2023). Typically, LLMs are augmented with an auxiliary audio encoder designed to preprocess audio signals, transforming them into the same input space as that of the LLMs, enabling them to achieve various speech tasks, such as automatic speech recognition (ASR), speech question answering (SQA), and so on. However, previous research has yet to overcome significant challenges in achieving effective generalization due to two main issues: (1) specialized tasks are highly sensitive to prompt design, resulting in performance degradation when confronted with unseen or complex instructions; (2) there is an absence of speech Chain-of-Thought (CoT) capability, which is essential for addressing complex tasks.
@@ 3 Method
Figure 1: Overview of the proposed WavLLM. Two speech encoders and adapters with different focuses are utilized, where Whisper and the corresponding adapter is used for extracting semantic information, and WavLM for extracting acoustic information. Before being fed to the LLM, these two representations are concatenated together and linearly transformed. Adaptive LoRA approach is used for cross-modal efficient fine-tuning with online adaptation, where the prompt adapter is able to generate prompt-dependent parameters to adjust the amplitude of LoRA in the second advanced multi-task training stage.
@@ 4.4 Analysis
Table 5 and 6 show the results of our models after first mixed single-task training stage and second advanced multi-task training stage. As discussed earlier, the model with only mixed training is incapable of generalizing to multi-task instructions, and the poor performance of II-tasks confirms this. After advanced training, our model produces comparable or even better performance on single-task prompts compared to the first-stage model. For zero-shot II-tasks, significant enhancement of generalization ability is obtained after advanced training, as evidenced not only by the increased adherence to instructions (92.50% vs. 26.25%) but also by the higher accuracy of each executed instruction (62.44% vs. 22.92%). For cross-lingual summary tasks using CoT based instructions, our model after the second advanced multi-task training stage consistently outperforms the first stage model. In addition, we found that the first stage model mainly accomplished the ST task and did not perform the summarization task. To better demonstrate the effectiveness of the second stage, we crafted a long story-based CoT task by GPT-4 where the audio contains a 100-word story, and the target is a 20-word summary in German. In this task, if the model solely focuses on completing the translation, there will be a noticeable discrepancy in length between its output and the desired target. From the results of this task in Table 6, the second advanced multi-task training stage model significantly outperforms the first stage model. Similar results are obtained when utilizing SALMONN-7B to conduct story-based CoT task instructions, which produces 10.6, 1.3, 7.8 and 63.90 on R-1, R-2, R-L and BERTScore respectively.

### v2

@@ 1 Introduction
Amidst the broadening scope of abilities, speech stands out as a crucial form of human communication, prompting extensive research to equip large language models (LLMs) with speech perception capabilities (Shu et al., 2023; Wu et al., 2023; Tang et al., 2023; Chu et al., 2023; Ma et al., 2024). Typically, LLMs are augmented with an auxiliary audio encoder designed to preprocess audio signals, transforming them into the same input space as that of the LLMs, enabling them to achieve various speech tasks, such as automatic speech recognition (ASR), speech question answering (SQA), and so on. However, previous research has yet to overcome significant challenges in achieving effective generalization due to two main issues: (1) specialized tasks are highly sensitive to prompt design, resulting in performance degradation when confronted with unseen or complex instructions; (2) there is an absence of speech Chain-of-Thought (CoT) capability, which is essential for addressing complex tasks.
@@ 3 Method
Figure 1: Overview of the proposed WavLLM. The left part (a) is the two-stage curriculum learning. The right part (b) is the model architecture. Two speech encoders and adapters with different focuses are utilized, where Whisper is used for extracting semantic information, and WavLM for extracting acoustic information. Before being fed to the LLM, these two representations are concatenated together and linearly transformed. Adaptive LoRA approach is used for cross-modal efficient fine-tuning with online adaptation, where the prompt adapter is able to generate prompt-dependent parameters to adjust the amplitude of LoRA in the second advanced multi-task training stage.
@@ 4.4 Analysis
Table H and 6 show the results of our models after first mixed single-task training stage and second advanced multi-task training stage. As discussed earlier, the model with only mixed training is incapable of generalizing to multi-task instructions, and the poor performance of II-tasks confirms this. After advanced training, our model produces comparable or even better performance on single-task prompts compared to the first-stage model. For zero-shot II-tasks, significant enhancement of generalization ability is obtained after advanced training, as evidenced not only by the increased adherence to instructions (92.50% vs. 26.25%) but also by the higher accuracy of each executed instruction (62.44% vs. 22.92%). For cross-lingual summary tasks using CoT based instructions, our model after the second advanced multi-task training stage consistently outperforms the first stage model. In addition, we found that the first stage model mainly accomplished the ST task and did not perform the summarization task. To better demonstrate the effectiveness of the second stage, we crafted a long story-based CoT task by GPT-4 where the audio contains a 100-word story, and the target is a 20-word summary in German. In this task, if the model solely focuses on completing the translation, there will be a noticeable discrepancy in length between its output and the desired target. From the results of this task in Table 6, the second advanced multi-task training stage model significantly outperforms the first stage model. Similar results are obtained when utilizing SALMONN-7B to conduct story-based CoT task instructions, which produces 10.6, 1.3, 7.8 and 63.90 on R-1, R-2, R-L and BERTScore respectively.
@@ Appendix E Example of Multi-round Dialog
Appendix F Training Data Details
Training data used in the first stage and second stage. For all tasks, the instructions are diverse. “#Hours” refers to the duration of speech data for each task, not the total number of hours of the data source. The targets of SQA tasks are generated using GPT3.5, GPT-4 or LLaMA-2-chat. Task Description Data Source #Hours Single -task automatic speech recognition (ASR) LibriSpeech 960 LibriHeavy medium 1800 speech-to-text translation (ST), including English to German (En2De), English to Japanese (En2Ja), and English to Chinese (En2Zh) CoVoST2 440 MuST-C 280 speaker verification (SV) VoxCeleb 1290 emotion recognition (ER) IEMOCAP Session 1-4 5 speech question answering (SQA), including gender and speaker-related questions, and multi-round QA LibriSpeech 520 AMI 50 Fisher 710 Switchboard 230 speech question answering (SQA), continue writing tasks LibriSpeech 960 speech question answering (SQA), summary tasks LibriSpeech 410 instruction tuning (IT), including spoken based and text based tasks Alpaca 90 Multi -task ER + text based IT IEMOCAP Session 1-4, Alpaca 71 ASR + text based IT LibriSpeech,  Alpaca 274 ST + text based IT CoVoST2, MuST-C, Alpaca 343 SV + text based IT VoxCeleb, Alpaca 243 SQA + text based IT AMI, Fisher,Switchboard, Alpaca 773 ASR + ST LibriSpeech 74 ASR + SQA LibriSpeech 43 ASR + ST + text-based IT CoVoST2, Alpaca 5 ASR + SQA + text-based IT LibriSpeech, Alpaca 1066
Appendix G Implementation Details
As mentioned above, the semantic and acoustic speech encoders are the encoder of Whisper-large-v2888https://huggingface.co/openai/whisper-large-v2 and WavLM-base [^9] https://huggingface.co/microsoft/wavlm-base, the backbone LLM is LLaMA-2-chat-7b [^10] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, and all of their parameters are frozen. The outputs of both modality adapters have a time stride of 80 ms and a dimension of 2048, and the rank ($R$) of LoRA is set as 32. In the first mixed single-task training stage, the total number of parameters in our model is 7.55 billion, of which 76.6 million are trainable. In the advanced training phase, the bottleneck dimension ($K$) of the prompt adapter is set as 1024. The 4096-dimensional prompt-dependent parameters produced by prompt adapter are element-wise multiplied with the outputs of the LoRA. Our models are trained with the two-stage curriculum-learning method on 32 V100 GPUs using the Adam optimizer, set with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.98$ and batch size equivalent to 30 seconds per GPU, where the first stage consisted of 400K steps and the subsequent stage involved an additional 150K steps. Additionally, we employed a maximum learning rate of $1\times 10^{-4}$, incorporating a warm-up phase for the first 10% of steps. The two-stage training data are presented in data construction part of Section 3.2.
Appendix H The Effect of Advanced Training for Single-tasks
Performance of model with or without advanced training on single-task instructions. mixed training means the first mixed single-task training stage, and advanced training means the second advanced multi-task training stage.
       Models ASR ST (En2De) SV ER SQQA SQA
test-clean test-other CoVoST2 MUSTC
WER BLEU Acc. Acc. Acc. Acc.
mixed training 2.0 4.8 23.9 21.9 0.91 0.72 0.55 67.30%
+ advanced training 2.0 4.8 23.6 21.7 0.91 0.72 0.57 67.55%